### v1

# LLM-Guided Open RAN: Empowering Hierarchical RAN Intelligent Control

Abstract
Recent advancements in large language models have led to a significant interest in deploying LLM-empowered algorithms for wireless communication networks. Meanwhile, open radio access network (O-RAN) techniques offer unprecedented flexibility, with the (non-RT RIC) and (near-RT RIC) components enabling intelligent resource management across different time scales. In this paper, we propose the LLM empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs. This framework integrates LLMs with reinforcement learning (RL) for efficient network resource management. In this framework, LLMs-empowered non-RT RICs provide strategic guidance and high-level policies based on environmental context. Concurrently, RL-empowered near-RT RICs perform low-latency tasks based on strategic guidance and local near-RT observation. We evaluate the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting. Simulation results demonstrate that the proposed framework achieves superior performance. Finally, we discuss the key future challenges in applying LLMs to O-RAN.
Index Terms:
Open radio access network, large language model, reinforcement learning
I Introduction
The O-RAN has recently gained significant attention for its ability to prompt interoperability and flexibility. In O-RAN, the next generation node B (gNB) is functionally split into three components: the radio unit (RU), the distributed unit (DU), and the central unit (CU). The different components of the networks can be developed and supplied by various vendors through open standards and interfaces. This openness enables a competitive ecosystem, driving innovation and reducing costs. Moreover, by introducing a modular, software-driven architecture, O-RAN transforms the traditional hardware-centric approach to virtual RAN [1].
A core component of O-RAN is the RICs, which provides advanced control and optimization capabilities. The RICs include: 1) the non-RT RIC, which integrates with the network orchestrator and operates on a time scale larger than 1 s, and 2) the near-RT RIC, which manages control loops with RAN nodes on a time scale between $10$ ms and $1$ s. Those RICs are capable of aggregating key performance measurements (KPMs) to monitor the status of network infrastructure (e.g., number of users, load, throughput, resource utilization). Specifically, the near-RT RIC connects to the RAN components through E2 interface. The non-RT RIC interfaces with RAN components through O1 interface, and based on the available KPMs, it facilitates a non-real-time control loop and supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via A1 interface. The third-party applications, known as xApp and rApp, are deployed within near-RT RIC and non-RT RIC, respectively. These applications enable advanced functionalities, such as enhanced radio resource management, improved energy efficiency, and predictive maintenance [2].
Significant researches have been conducted to enhance RICs functionalities [3, 4, 5, 6, 7, 8]. Specifically, near-RT RIC, equipped with xApps, manages a smaller number of devices and enable low-latency optimizations. These optimizations include radio card switching to improve energy efficiency [3], flow-split distribution, congestion control, and scheduling to optimize network utility [4]. In contrast, the non-RT RIC, equipped with rApps, is responsible for managing a larger number of devices and performing high-level, long-term network management [5, 6, 7, 8]. For example, the orchestRAN framework in [5] automatically computes the optimal set of data-driven algorithms and determines their execution locations. Furthermore, cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides near-RT RIC with high-level guidance, including network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then leverages real-time information to refine and implement these policies, enhancing dynamic network performance. Despite these advancements, two critical challenges remain: 1) Traditional machine learning (ML) techniques struggle to process the diverse and complex information in O-RAN. 2) Ensuring high-quality policy generation in the non-RT RIC across diverse scenarios and improving efficient cooperation between non-RT RIC and near-RT RIC are challenging.
Meanwhile, the LLMs, such as OpenAI’s ChatGPT, Google’s Gemini, and Meta AI’s Llama, are advanced ML models, designed to understand and generate human language. Trained on vast and diverse datasets, LLMs excel in tasks such as text generation, translation, summarization, and question answering. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Telecom operators can leverage LLMs to enhance network management and improve overall service quality. Additionally, LLMs support the development of intelligent applications for network operations, including predictive maintenance, anomaly detection, and performance optimization. The LLMs have opened new routes for wireless network design and operation, offering the potential for significant performance improvements.
Recent studies have applied LLM-based technologies to wireless network for resource allocation[9, 10, 11], network deployment[12, 13] and packet analysis [14]. Specifically, [9] develope a LLM empowered-xApp to allocate physical resource block (PRB) for 5G O-RAN. However it lacks a comparison with state of art resource allocation algorithms and does not analyze the inference latency of LLMs. [10] and [11] utilize Meta AI’s Llama to derive the power allocation policy directly. Additionally, [12] deploys a combined LLM and conventional neural network (CNN) model as a RL agent for base station placement. Furthermore, the GPT model has been utilized to formulate mathematical models and generate solution codes for base station placement [13]. Although these applications show the potential of LLMs in wireless networks, several challenges remain. First, the high computational complexity hinders near-RT decision-making. Additionally, without fine-tuning for domain-specific tasks, their performance often remains comparable to traditional algorithms.
Motivated by these opportunities and challenges, we propose the LLM-hRIC framework for O-RAN. This framework deploys LLMs-empowered non-RT RIC to provide guidance for the RL-empowered near-RT RIC. The near-RT RIC utilizes this guidance along with near-RT environment information to perform near-RT decision-making. The contributions of this work are summarized as follows:
  • LLM-empowered non-RT RIC: We propose a LLM-empowered non-RT RIC that collects and processes comprehensive environmental data to provide strategic guidance to the near-RT RIC. This approach improves data integration and long-term decision-making in the O-RAN framework.
  • RL-empowered near-RT RIC: We introduce a RL-empowered near-RT RIC that utilizes the guidance from non-RT RIC and local observations to make near-RT decisions. To accelerate convergence and improve overall system performance, we deploy a joint training method that smoothly mixes non-RT RIC’s guidance and near-RT RIC’s policy.
  • Performance Evaluation: We evaluate the proposed framework in IAB networks. The results show significant performance improvements compared to baseline methods, demonstrating its ability to manage the complex information in O-RAN systems.
II LLM-empowerd hierarchical RIC Framework
In this section, we propose the LLM-empowerd hierarchical RIC (LLM-hRIC) framework for the resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages on generalization across various environments with minimal language data, it is used for the non-RT RIC to provide a guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Proposed LLM-hRIC framework.
II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, is responsible for utilizing those information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enables to form a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide a high-quality strategic guidance to the near-RT RIC.
The process of the LLM-empowered non-RT RIC can be summarized as follows.
  • 1)
    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images), and integrates the collected data by structuring it into an unified format.
  • 2)
    Environment Understanding and Actionable Insights Generation: The LLM analyze and understand the context based on the integrated multimodal data, and generate a strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control). The mathematical formulation of the guidance generation process is given as
    $\textbf{y}=\text{LLM}_{\theta}(x_{1},x_{2},\cdots,x_{N}),$ (1)
    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$, and $\theta$ denotes the parameters of LLM.
  • 3)
    Guidance transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via A1 interface to be used for the near-RT decision-making process.
II-B RL-Empowered near-RT RIC
The near-RT RIC operates as a politic layer of O-RAN, managing and optimizing network operations in near-RT. To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface. Specifically, the optimal action $\boldsymbol{a}$ can be determined as
$\boldsymbol{a}=\text{RL}_{\phi}(\textbf{y},\boldsymbol{o}),$ (2)
where $\phi$ represents the parameters of the RL model.
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the RL-based xApps need to be designed to process information efficiently and make a decision fast. From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments. This can enable the RL-based xApps to make a decision with the large-scale view on the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
III Use Case: Power Allocation in IAB Network
In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework.
III-A Power Allocation for IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as =$\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users, $\mathcal{K}$, follow a Gaussian-Markov mobility model and always connect to the nearest SBS. The SBS allocates fixed power to their connected user.
The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. The large-scale fading gain of LoS and NLoS components are given by $L^{\text{LoS}}$ and $L^{\text{NLoS}}$, respectively. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
In this use case, the total channel bandwidth $W$ for each MBS is allocated between the backhaul and access links in proportion $\alpha$, where $\alpha\in[0,1]$, i.e., the fraction of bandwidth allocated to the backhaul link is $W\alpha$ and the remaining bandwidth is assigned to the access links. At the $t$-th time slot, the achievable rate of the access link between the $k$-th user and the $n$-th SBS and that of the backhaul link between the $m$-th MBS and the $n$-th SBS are denoted by $R_{m,n,k}^{\text{a}}[t]$ and $R_{m,n}^{\text{b}}[t]$, respectively, and then the total throughput becomes $\mathcal{T}=\sum_{t=1}^{T}\sum_{m=1}^{M}\sum_{n\in\mathcal{N}_{m}}\min\left(R_{m,n}^{\text{b}}[t],\sum_{k\in\mathcal{K}_{m,n}}R_{m,n,k}^{\text{a}}[t]\right)$. The objective is to maximize the total throughput $\mathcal{T}$ by optimizing the power allocation of MBSs to each SBSs under the total power constraint on the MBS, given by $\sum_{n\in\mathcal{N}_{m}}P_{m,n}[t]\leq P_{m}^{\text{max}},\forall m,t$, where $P_{m}^{\text{max}}$ is the maximum available power of the $m$-th MBS.
III-B LLM-hRIC Framework for Power Allocation
In this section, the proposed LLM-hRIC framework is deployed for power allocation in the IAB networks. Each MBS connects to a near-RT RIC, while a centralized non-RT RIC collects information from all MBSs and provides strategic power allocation guidance to the near-RT RICs.
III-B1 LLM-Empowered non-RT RIC
In this framework, the LLM is deployed in the non-RT RIC as a rApp. A structured prompt guides the LLM in generating the power allocation guidance for near-RT RIC (e.g., initial power allocation policy). This prompt includes the following components:
  • Objective Function: Defines the goal of power allocation optimization, i.e., to maximize the total throughput.
  • Environment Description: Details the number of MBSs, SBSs, the maximum power of each MBS, etc,.
  • Channel Information: Provides the average channel gain between MBSs and SBSs.
  • Interference Information: Specifies interference sources and their average interference channel gains.
  • Constraints: Constraints for power allocation policy and the guidance format.
The structure of the prompt is as follows:
Prompt Structure You are an expert in wireless communications for resource allocation. Your objective is <Objective function >. Here is the system description: <Environment information >. Input Format: For each MBS, you are provided with the following input data for its connected SBSs. Each SBS is represented as a list: [Average channel gain, number of connected users, average expected data rate of connected users (Mb/s), and interference: [interference source (MBS,SBS), average interference channel gain] ]. Constraints: Ensure the total power allocation across SBSs for each MBS sums to 1. For each MBS, output the normalized power allocation ratios as a list of six values corresponding to its six SBSs. MBSX: [value1, value2, value3, value4, value5, value6].
Due to the time-varying nature of small-scale fading, the non-RT RIC uses average values of small-scale fading parameters to compute metrics like channel gain, expected data rate, and interference. These average values effectively represent the channel’s state in future time slots. The LLM generates the initial power allocation policy as a guidance, denoted as $\boldsymbol{p}^{o}_{m}$, for the $m$-th MBS. This guidance is updated every $T_{0}$ time slots to adapt to network environment change.
III-B2 RL-Empowered near-RT RIC
In the IAB networks, each MBS is associated RL-based xApps for near-RT resource management. The markov decision process (MDP) modeled as following components:
  • State Space: We define the system state space of $m$-th MBS as $\mathcal{S}_{m}=\{\boldsymbol{h}_{m},\boldsymbol{n}_{m},\boldsymbol{R}_{m},\boldsymbol{p}^{o}_{m}\}$, where $\boldsymbol{h}_{m}$ is the channel gain between MBS and its connected SBSs, $\boldsymbol{n}_{m}$ represents the number of associated users per connected SBS, $\boldsymbol{R}_{m}$ is the average expected data rate of connected users per SBS, and $\boldsymbol{p}^{o}_{m}$ is the initial power allocation policy provided by the non-RT RIC.
  • Action Space: Each MBS allocates power to its connected SBSs, denoted as $\boldsymbol{p}_{m}=(p_{m,1},\cdots,p_{m,N})$, where $p_{m,n}\in[0,1]$ is the power allocation ratio for each connected SBS.
  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS, and $r^{\text{g}}$ is the total throughput of all MBSs.
III-C Cooperative Training
The RL agent in the near-RT RIC need to be trained to effectively utilize the guidance provide by LLM. Traditional RL typically relies on random noise for action exploration during the initial training stages. To enhance exploration efficiency, the initial actions are instead guided by the LLM, e.g., initial power allocation policy. And then the RL action will mix with LLM guidance gradually. The training is divided into the following three phases:
  • Phase 1: The power allocation policy $\boldsymbol{p}_{m}$ is generated by adding noise to the policy $\boldsymbol{p}_{m}^{o}$ provided by the LLM, facilitating exploration around the LLM’s recommendations.
  • Phase 2: The power allocation policy $\boldsymbol{p}_{m}$ is obtained by linearly blending the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$ using a decay weighted $w$: $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where $w$ gradually decays from $1$ to 00 during this phase.
  • Phase 3: The power allocation policy $\boldsymbol{p}_{m}$ is determined by the RL agent.
In each training stage, the power allocation policy interact with environment, and the resulting MDP experiences are collected to continuously train the RL agent.
III-D Experimental Results
In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 1s). Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 100ms). To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 2: Training curve for all methods when $\alpha=0.5$.
Figure 2 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$. The results show that the total throughput increases steadily as training progresses and gradually converges as the models learn effective policies. Notably, the proposed LLM-hRIC framework achieves higher total throughput compared to the baseline methods. This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance.
Figure 3: Effect of backhaul-access bandwidth partitioning ratio $\alpha$ on total throughput.
Figure 3 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput. The result shows that the throughput initially increases and then decreases as the allocation ratio $\alpha$ increases. This is because the performance bottleneck exists at the backhaul link for a low $\alpha$, while it is primarily in the access link for large $\alpha$. The proposed LLM-hRIC framework consistently outperforms baseline methods. This is due to the high-quality initial policies, generated by LLM, which enhances exploration efficiency and facilitates better policy discovery by RL. The LLM leverage detailed, and global information to improve coordination among MBSs, resulting in a more efficient resource allocation and higher overall network performance.
IV Future Challenges
In this section, we discusses the future challenges of LLMs in O-RAN. As shown in Fig. 4, key challenges include leveraging multi-modal data, fine-tuning of LLMs, enabling systematic collaboration between RICs, and achieving timely inference. Each topic is discussed in detailed below.
Figure 4: Future challenges.
IV-A Leveraging Multi-Model Data
Multi-modal data (e.g. text, signals, images, and videos) can be used for better understanding on the complex network environments. The multi-modal LLMs are capable of reasoning across various data types, providing richer contextual insights and enabling more informed decision-making.
Meanwhile, the zero-shot or few shot learning methods (e.g. retrieval augmentd generation (RAG) and prompt engineering) have been explored to adapt LLMs to new environments. However, several open questions remain. First, current prompt engineering methods generally lack support for multi-modal inputs in wireless communication domain, making it critical to develop methods that effectively handle such data. Moreover, the performance of RAG is highly depend on retrieval engine. Designing an effective multi-modal RAG engine in wireless communication domain is key challenge for improving LLM reasoning.
IV-B Fine-Tuning of LLMs
While current LLMs have been trained well for general domain, they are not specifically optimized for wireless communication tasks, which limits their performance in this area. Although, prompt engineering can guide LLMs to some extent, limitations in context window size and the dynamic nature of wireless environment hinder real time reasoning.
To address this issue, fine-tuning LLMs using offline, domain-specific data is essential. Fine-tuning LLMs can enhance decision-making relevance and reliability. Nonetheless, two major open challenges remain: 1) how to collect or effectively utilize limited wireless communication data, and 2) how to effectively generate high quality instruction data for fine-tuning?
IV-C Systematic Collaboration between RICs
Due to the lack of domain specific fine-tuning, LLMs-generated policies are not always optimal. Enhancing them with lightweight AI techniques is critical, especially for near-RT applications where inference speed is paramount. To fully exploit the strength of both LLMs (for long time, global optimization) and lightweight AI (for real-time responsiveness), a cooperative framework must be established. This framework should enable smooth interaction between the LLMs and lightweight AI models, supporting adaptive and robust decision-making under dynamic conditions.
Several challenges need to be addressed. First, LLMs operate on large time scale, while network environment change rapidly in near-RT. Bridging this time gap to make training and prediction work well is not easy. Second, due to the limited capacity of lightweight models, selecting the most relevant features for specific tasks is critical for performance enhancement.
IV-D Timely Inference
Modern LLMs, such as GPT-4 with 1.7 trillion parameters, demand significant computational resources and are often deployed on cloud-based servers. This dependency introduces substantial latency, making deploying LLM model in near-RT RIC as a xApp challenging.
To address this challenge, technologies such as model pruning, quantization and knowledge distillation are crucial for reduce the model size. Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead without significantly affecting performance. In addition, the model pruning and knowledge distillation train a smaller, optimized versions of LLMs. The different methods can compress the model with different ratio. However Compression methods typically trade off some model performance for reduced latency. Balancing model size, inference speed, and accuracy is a critical challenge.
V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM is deployed as rApp in the non-RT RIC to provide strategic guidance, while RL is implemented as xApp in the near-RT RIC to facilitate near-RT decision-making. The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems.