### v3

# arXiv:2504.18062v3[cs.NI] 27 Feb 2026

Abstract
Despite recent advances in applying large language models and machine learning (ML) techniques to open radio access network (O-RAN), critical challenges remain, such as insufficient cooperation between RICs, high computational demands hindering real-time decisions, and the lack of domain-specific fine-tuning. Therefore, this article introduces the LLM-empowered hierarchical RIC (LLM-hRIC) framework to improve the collaboration between RICs in O-RAN. The LLM-empowered non-real-time (non-RT RIC) acts as a guider, offering a strategic guidance to the near-real-time (near-RT RIC) using global network information. The RL-empowered near-RT RIC acts as an implementer, combining this guidance with local real-time data to make near-RT decisions. We evaluate the feasibility and performance of the LLM-hRIC framework in an integrated access and backhaul (IAB) network setting, and finally, discuss the open challenges of the LLM-hRIC framework for O-RAN.
I Introduction
The O-RAN has recently gained significant attention for its ability to prompt interoperability and flexibility. In O-RAN, the next generation node B (gNB) is functionally split into three components: the radio unit (RU), the distributed unit (DU), and the central unit (CU). The different components of the networks can be developed and supplied by various vendors through open standards and interfaces. This openness enables a competitive ecosystem, driving innovation and reducing costs. Moreover, by introducing a modular, software-driven architecture, O-RAN transforms the traditional hardware-centric approach to virtual RAN [1].
A core component of O-RAN is the RICs, which provides advanced control and optimization capabilities. The RICs include: 1) the non-RT RIC, which integrates with the network orchestrator and operates on a time scale larger than 1 s, and 2) the near-RT RIC, which manages control loops with RAN nodes on a time scale between $10$ ms and $1$ s. The non-RT RIC and the near-RT RIC are capable of obtaining key performance measurements (KPMs) via O1 and E2 interfaces, respectively, to monitor the status of network infrastructure (e.g., number of users, load, throughput, and resource utilization). The non-RT RIC also supports the deployment of policies, guidance, and intelligent models within the near-RT RIC via the A1 interface. Specific RIC functionalities, such as mobility management and network slicing, are deployed within near-RT RIC and non-RT RIC in the form of the third-party applications, known as xApp and rApp, respectively [2].
Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8, 9]. Specifically, the near-RT RICs has been developed to manage a smaller number of devices and optimize on a small-time scale, such as the radio card switching for energy efficiency [3] and the flow-split distribution, congestion control, and scheduling for higher network utility [4]. In particular, deep reinforcement learning (DRL) has emerged as a key technology for intelligent xApps, enabling dynamic control of radio resource management in O-RAN [9]. The non-RT RIC has been developed to manage a larger number of devices and perform high-level and long-term network management, such as the orchestRAN[5], which automatically computes the optimal set of algorithms and their execution location. Furthermore, the cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides the high-level guidance to near-RT RIC, including the network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then refines and implements these policies based on real-time information to enhance the network performance.
Meanwhile, the LLMs have opened new routes for wireless network design and operation including O-RAN. The LLMs are advanced ML models, designed to understand and generate human language. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Recent studies have applied LLM-based technologies to wireless network for the resource allocation[10, 11], base station placement[12], and packet analysis [13]. Recently, the LLM-based xApp to allocate physical resource block (PRB) for O-RAN has also been introduced[14].
Despite advancements in RICs and LLM-based techniques for wireless networks, a critical real-world problem remains: bridging the ‘coordination gap’ between the strategic, long-term control of the non-RT RIC and the tactical, real-time execution of the near-RT RIC. This gap is exacerbated by several underlying challenges: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
Motivated by these opportunities and challenges, we propose the novel framework for O-RAN, which is the LLM-hRIC. In this framework, there are two RICs: 1) the LLMs-empowered non-RT RIC and 2) the reinforcement learning (RL)-empowered near-RT RIC. The LLMs-empowered non-RT RIC works as a guider to the near-RT RIC. Based on the objective of the near-RT RIC, it provides strategic guidance for the RL-empowered near-RT RIC based on the global view of the networks (e.g., long-term channel information, network nodes state, users mobility model, and users distribution). The RL-empowered near-RT RIC works as an implementer, which utilizes this guidance along with the local view of the networks (e.g., real-time channel quality and local network node state) to perform near-RT decision-making. The key contribution of the paper can be summarized as follows.
  • We propose the novel LLM-hRIC framework for O-RAN, which can show how the LLM can be utilized to enhance the performance of RICs in O-RAN. This article focuses on addressing the limitation of LLM in making near-RT decisions and paves a way to utilize it in a reliable way in O-RAN environments.
  • We introduce the three sequential phases for training of the RL-empowered near-RT RIC, of each phase utilizes the guidance of the non-RT RIC in a different way to accelerate convergence as well as improve the system performance.
  • We quantify the performance of the LLM-hRIC for the power allocation in the IAB networks as a use case. This can demonstrate the ability of LLM-hRIC to manage the complex information in O-RAN.
II LLM-empowerd hierarchical RIC Framework
In this section, we introduce the LLM-empowered hierarchical RIC (LLM-hRIC) framework for resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages of generalization across various environments with minimal language data, it is used for the non-RT RIC to provide guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Structure of the proposed LLM-hRIC framework and time diagrame.
II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, are responsible for utilizing that information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multi-modal data to form a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide high-quality strategic guidance to the near-RT RIC.
The process of the LLM-empowered non-RT RIC can be summarized as following steps.
  • 1)
    Data Integration: The non-RT RIC ingests multi-modal data from gNB and near-RT RIC, primarily via the O1 interface. These data include long-term time-series metrics (e.g., total throughput trends), unstructured textual data (e.g., fault alarm logs), topology data representing network node connectivity, and even advanced visual data such as radio maps generated from digital twins.
  • 2)
    Input Validation and Prompt Construction: Before being formatted into a prompt, this data undergoes a validation stage to handle potentially noisy or corrupted measurements. The data integration process involves structuring these varied inputs into a unified, context-rich format that can be effectively processed by the LLM.
  • 3)
    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate the text response, that includes the strategic guidance y (e.g., transmit power allocation strategies).
  • 4)
    Guidance Extraction and Verification: The LLM’s raw text output is validated for its conformance to predefined formats and constraints, and extracts the strategic guidance y. To ensure robustness, if this process fails, a fallback mechanism provides a default policy as guidance.
  • 5)
    Guidance Transfer to the near-RT RIC: The guidance vector y is then sent to the near-RT RIC via the A1 interface, used for the near-RT decision-making process.
II-B RL-Empowered near-RT RIC
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the xApps need to be designed to process information efficiently and make fast decisions in a highly dynamic network environment. Hence, we employ RL because its ability to learn an adaptive policy is well-suited for reacting to the dynamic environments with real-time changes. Furthermore, RL is naturally designed to incorporate a strategic guidance vector as a part of its state input, fusing it with the observed local information to make refined and context-aware decisions.
The process of the RL-empowered near-RT RIC can be summarized in the following steps:
  • 6)
    Data Integration: The near-RT RIC ingests the strategic guidance from the non-RT RIC (via the A1 interface) and the observed local information from the gNB (via the E2 interface).
  • 7)
    Policy generation: The RL-based xApp then analyzes these inputs to generate an optimal, fine-grained policy that is sent to the gNB for execution.
This integration of global insights with real-time local information enhances the RL agent’s ability to make refined, fine-grained decisions that are both strategically sound and tactically optimal for the immediate radio conditions.
III Use Case: Power Allocation in IAB Network
In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework. This scenario was chosen as a foundational proof-of-concept since the resource allocation in multi-cell IAB networks is a well-known difficult problem to handle as it requires the network-wise global information and control. This problem can demonstrate our framework’s strength: using the LLM in the non-RT RIC to provide a globally-aware and coordinated guidance to local RL agents.
III-A Power Allocation in IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as $\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users $\mathcal{K}_{m,n}$, connected with the $n$-th SBS of the $m$-th MBS, follow a Gaussian-Markov mobility model. The SBS allocates fixed power to their connected user. The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
In this use case, the total channel bandwidth $W$ for each MBS is allocated between the backhaul and access links in proportion $\alpha$, where $\alpha\in[0,1]$, i.e., the fraction of bandwidth allocated to the backhaul link is $W\alpha$ and the remaining bandwidth is assigned to the access links. At the $t$-th time slot, the achievable rate of the access link between the $k$-th user and the $n$-th SBS and that of the backhaul link between the $m$-th MBS and the $n$-th SBS are denoted by $R_{m,n,k}^{\text{a}}[t]$ and $R_{m,n}^{\text{b}}[t]$, respectively, and then the total throughput becomes $\mathcal{T}=\sum_{t=1}^{T}\sum_{m=1}^{M}\sum_{n\in\mathcal{N}_{m}}\min\left(R_{m,n}^{\text{b}}[t],\sum_{k\in\mathcal{K}_{m,n}}R_{m,n,k}^{\text{a}}[t]\right)$. The objective is to maximize the total throughput $\mathcal{T}$ by optimizing the power allocation of MBSs to each SBSs under the total power constraint on the MBS, given by $\sum_{n\in\mathcal{N}_{m}}P_{m,n}[t]\leq P_{m}^{\text{max}},\forall m,t$, where $P_{m}^{\text{max}}$ is the maximum available power of the $m$-th MBS.
III-B LLM-hRIC Framework for Power Allocation
In this section, the proposed LLM-hRIC framework is deployed for power allocation in the IAB networks. Each MBS connects to a near-RT RIC, while a centralized non-RT RIC collects information from all MBSs and provides strategic power allocation guidance to the near-RT RICs.
III-B1 LLM-Empowered non-RT RIC
In this framework, the LLM is deployed in the non-RT RIC as a rApp. A structured prompt guides the LLM in generating the power allocation guidance for near-RT RIC (e.g., initial power allocation policy). This prompt includes the following components:
  • Objective Function: Defines the goal of power allocation optimization, i.e., to maximize the total throughput.
  • Environment Description: Details the number of MBSs, SBSs, the maximum power of each MBS, etc,.
  • Channel Information: Provides the average channel gain between MBSs and SBSs.
  • Interference Information: Specifies interference sources and their average interference channel gains.
  • Constraints: Details the constraints on power allocation policy and the guidance format.
The structure of the prompt is as Fig. 2.
Prompt Structure You are an expert in wireless communications for resource allocation. Your objective is <Objective function >. Here is the system description: <Environment information >. Input Format: For each MBS, you are provided with the following input data for its connected SBSs. Each SBS is represented as a list: [Average channel gain, number of connected users, average expected data rate of connected users (Mb/s), and interference: [interference source (MBS, SBS), average interference channel gain] ]. Constraints: Ensure the total power allocation across SBSs for each MBS sums to 1. For each MBS, output the normalized power allocation ratios as a list of six values corresponding to its six SBSs. MBSX: [value1, value2, value3, value4, value5, value6].
Figure 2: The Prompt Structure.
Due to the time-varying nature of small-scale fading, the non-RT RIC uses average values of small-scale fading parameters to compute metrics like channel gain, expected data rate, and interference. These average values effectively represent the channel’s state in future time slots.
To ensure the reliability of these inputs, a data checking step performs range validation. This process verifies that all metrics are physically plausible, such as ensuring that the ‘number of connected users’ is non-negative and that the ‘average channel gain’ is within a realistic range.
III-B2 RL-Empowered near-RT RIC
In the IAB networks, each MBS is associated RL-based xApps for near-RT resource management. The markov decision process (MDP) modeled as following components:
  • State Space: We define the system state space of $m$-th MBS as $\mathcal{S}_{m}=\{\boldsymbol{h}_{m},\boldsymbol{n}_{m},\boldsymbol{R}_{m},\boldsymbol{p}^{o}_{m}\}$, where $\boldsymbol{h}_{m}$ is the channel gain between MBS and its connected SBSs, $\boldsymbol{n}_{m}$ represents the number of associated users per connected SBS, $\boldsymbol{R}_{m}$ is the average expected data rate of connected users per SBS, and $\boldsymbol{p}^{o}_{m}$ is the initial power allocation policy provided by the non-RT RIC.
  • Action Space: Each MBS allocates power to its connected SBSs, denoted as $\boldsymbol{p}_{m}=(p_{m,1},\cdots,p_{m,N})$, where $p_{m,n}\in[0,1]$ is the power allocation ratio for each connected SBS.
  • Reward Function: The reward is designed to encourage individual performance and global cooperation $r_{m}=r^{\text{l}}_{m}+r^{\text{g}}$, where $r^{\text{l}}_{m}$ is the throughput of $m$-th MBS and $r^{\text{g}}$ is the total throughput of all MBSs.
III-C LLM-assisted Three Phases Training for RL
The RL agent in the near-RT RIC must be trained to effectively utilize the guidance provided by LLM. Traditional RL typically relies on a random policy for action exploration during the initial stage of training. To improve the exploration efficiency, this random policy is replaced with the LLM guidance, such as an initial power allocation policy and injected action noise. The policy is gradually formed by blending the actions, generated by the RL agent with the LLM guidance, until the agent is capable of making decisions independently. Specifically, the training can be processed by the following three sequential phases:
  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$. This encourages exploration in the vicinity of the LLM’s guidance, allowing the RL agent to collect high-quality experiences and learn a near-optimal initial policy.
  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 0 over time. This creates a smooth transition from guided exploration to self-directed optimization.
  • Phase 3 (Self-directed Decision): The power allocation policy $\boldsymbol{p}_{m}$ is determined solely by the RL agent, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{d}$.
In each training stage, the power allocation policy interacts with the environment, and the resulting MDP experiences are collected to train the RL agent. This phased approach prevents the instability seen with abrupt policy changes, allowing the agent to efficiently fine-tune its near-optimal policy, which leads to stable convergence and the highest overall throughput.
III-D Experimental Results
Our simulation deploys the Llama-3.1-8B-instruct model [15] in the rApp to generate a power allocation guidance policy on a 2s timescale. For efficient deployment, the model is loaded with bfloat16 precision, with inference parameters of temperature =0.6=0.6 and top_p =0.9=0.9. Meanwhile, a deep deterministic policy gradient (DDPG) agent in the xApp provides a fine-grained power allocation policy for a MBS on a 200ms timescale. The DDPG agent was trained using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 256.  [^1] Note that the demonstration on a prototype testbed can be done once O1 and A1 interfaces become sufficiently mature in open-source O-RAN platforms. To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 3: LLM and DDPG inference latency with 500 examples.
Fig. 3 shows the LLM and DDPG inference latency with $500$ examples. The LLM inference latency is less than 1.5 seconds, confirming its decision time is well within the operational budget required for the LLM-hRIC ($\leq$ 2s). Moreover, the DDPG inference time is less than $0.07$ ms. This extremely low latency is significantly faster than the 10 ms requirement for the most demanding near-real-time control.
Figure 4: Training curve for proposed with different blending coefficient $w$ and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
Fig. 4 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$. The results show that the total throughput increases steadily as training progresses and gradually converges as the models learn effective policies. Notably, the proposed LLM-hRIC framework achieves higher total throughput compared to the baseline methods. This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance. Furthermore, the decaying $w$ is critical for performance. A fixed $w=0$ leads to training failure from an abrupt policy shift, while a fixed $w=0.9$ causes performance to stagnate by suppressing the agent’s exploration. In contrast, the proposed decaying $w$ ensures a smooth transition from guided exploration to self-directed optimization, which is necessary for stable convergence to the highest-performing policy.
Figure 5: Average total throughput of the proposed LLM-hRIC and baseline methods according to backhaul-access bandwidth partitioning ratio $\alpha$ over 50 test examples.
Fig. 5 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput. The result shows that the throughput initially increases and then decreases as the allocation ratio $\alpha$ increases. This is because the performance bottleneck exists at the backhaul link for a low $\alpha$, while it is primarily in the access link for large $\alpha$. The proposed LLM-hRIC framework consistently outperforms baseline methods. This is due to the high-quality initial policies, generated by LLM, which enhances exploration efficiency and facilitates better policy discovery by RL. The LLM leverage detailed, and global information to improve coordination among MBSs, resulting in a more efficient resource allocation and higher overall network performance.
IV Future Challenges
In this section, we discusses the future challenges of LLM-hRIC for O-RAN. As shown in Fig. 6, key challenges specific to LLM-hRIC include leveraging multi-modal data, fine-tuning LLMs for guidance generation, enabling systematic collaboration between RICs in LLM-hRIC, and achieving timely inference. Each topic is discussed in detailed below.
Figure 6: Future challenges of the LLM-hRIC for O-RAN.
IV-A Leveraging Multi-Model Data in LLM-hRIC
In the LLM-hRIC, the multi-modal LLMs can reason across various data types. Multi-modal data (e.g., text, signals, images, and videos) provides detailed environment insights, enhancing the ability of the non-RT RIC to generate more accurate and context-aware guidance for near-RT RIC. Moreover, zero-shot and few-shot learning techniques, such as prompt engineering and retrieval augmentd generation (RAG), have been explored to enhance the performance of LLMs in new environments. Despite their potential, significant challenges remain: 1) prompt engineering lacks robust support for multi-modal inputs in the wireless communication domain, necessitating new methods to handle such data for generating high-quality guidance. 2) RAG performance relies heavily on the retrieval engine. Designing an effective multi-modal RAG engine for LLM-hRIC in the wireless communication domain is critical to improve LLM reasoning and guidance generation.
IV-B Fine-Tuning of LLMs
Current LLMs are trained for general domains but not optimized for wireless communication tasks, limiting guidance quality in LLM-hRIC. While prompt engineering can partially bridge this gap, its effectiveness is limited by the quality of examples and input token window size. Moreover, the guidance from a general-purpose LLM can sometimes be unstable, making fine-tuning important for adapting to environmental changes and ensuring reliability. Fine-tuning LLMs with offline, domain-specific data is essential to enhance guidance reliability in the non-RT RIC. However, three challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging. 2) Generating high-quality instruction data for fine-tuning is essential. Automated or semi-automated methods for generating diverse, domain-relevant instruction data must be developed to improve the guidance quality of fine-tuned LLMs. 3) A privacy-preserving architecture should be considered such as the combination of federated learning (FL) with differential privacy (DP). The DP adds controlled noise to the shared model updates in FL for a mathematical guarantee that private information on the training data is not leaked.
IV-C Systematic Collaboration between RICs
In LLM-hRIC, the LLMs-empowered non-RT RIC generates guidance for multiple near-RT RICs, each responsible for diverse tasks. A systematic collaboration framework is needed to ensure smooth interaction between the LLMs and lightweight AI models. Several challenges need to be addressed. 1) non-RT RICs operates on a large time scale, while near-RT RICs can operate on different small time scales. This temporal mismatch complicates the synchronization between the guidance generation in the non-RT RIC and the real-time decision-making in the near-RT RIC. 2) near-RT RICs perform a variety of tasks (e.g., power allocation, PRB allocation, and user association). Each task requires specific guidance tailored to its unique requirements, making it crucial to develop a mechanism for providing task-aware, customized guidance for different near-RT RICs.
IV-D Timely Inference
Modern LLMs, such as GPT-4 with 1.7 trillion parameters, require significant computational resources and are often cloud-deployed, causing latency issues. This makes deploying an LLM in the non-RT RIC challenging, as some non-RT RIC require latency near 1s. Technologies like model pruning, quantization, and knowledge distillation can reduce model size. Specifically, the quantization technology reduces the precision of model parameters to lower computational overhead, while pruning and distillation create smaller, optimized LLMs. These methods compress models at varying ratios but often trade off performance for speed. Balancing model size, inference speed, and accuracy is critical to ensuring efficient deployment of LLMs in the LLM-hRIC framework.
V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores the guidance with local observation to facilitate near-RT decision-making. The superior performance demonstrates the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems. For instance, the framework is extensible to advanced paradigms like the intent-based networking, where the LLM can translate an administrator’s high-level and multi-objective intent into strategic guidance. This guidance is then executed by RL agents with real-time and tactical adaptations, showcasing a powerful synergy for future network automation. Additionally, we identify key challenges for LLM-hRIC, as discussed in Section ‘Future Challenges’, to guide future research and development.